OpenStudy (anonymous):

Find two numbers with a sum of 20 and a difference of 14

OpenStudy (anonymous):

17 and 3.

OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) –3 –1 0 2 3 5 6 8

OpenStudy (anonymous):

what is the linear function?

OpenStudy (johnweldon1993):

if you want to do it out...set up the 2 equations x + y = 20 x - y = 14 solve for y in the first one y = 20 - x plug this into the second equation wherever you see a "y" x - (20-x) = 14 distribute that "-" sign x - 20 + x = 14 add 20 to both sides x + x = 34 2x = 34 x = 17 go back and plug that into your first equation to find y 17 + y = 20 y = 3

OpenStudy (anonymous):

^ i didnt even think about that @johnweldon1993 , i just was going through it in my head.

OpenStudy (anonymous):

A. f(x) = x + 2 B. f(x) = 3x C. f(x) = x – 2 D. f(x) = x – 1

OpenStudy (johnweldon1993):

x f(x) -3 -1 0 2 3 5 6 8 i believe this is what he has for the linear function part

OpenStudy (anonymous):

plug in each option.

OpenStudy (johnweldon1993):

@Hope_nicole yeah i know me too..but i wanted to do it out...so he could see how to solve any other problem like that

OpenStudy (anonymous):

for example, on the first one... f(x)=x+2 plug in the first value of -3 f(x)=-3+2 f(x)=-1... therefore the first set it right, and keep going until you find an equation that satisfies all the points.

OpenStudy (johnweldon1993):

well lets see... when x = -3....it adds what to get 1? 2 x = 0 ...it adds what to get 2? 2 x = 3...it adds what to get 5? 2 etc... so it looks like it adds 2 to whatever x is going to be....so that is?

OpenStudy (anonymous):

so it's b

OpenStudy (anonymous):

no check again @jumpman97

OpenStudy (anonymous):

The scatter plot shows the number of students per class at Monida Middle School and the number of magazine subscriptions each class sold for a fund-raiser. About how many subscriptions did the class of 30 students sel

OpenStudy (anonymous):

scattered plot please

OpenStudy (anonymous):

A. about 185 subscriptions B. about 175 subscriptions C. about 165 subscriptions D. about 195 subscriptions

OpenStudy (anonymous):

1 –7 2 –10 3 –13 4 –16

OpenStudy (anonymous):

what are the equations?

OpenStudy (johnweldon1993):

you cant figure it out @Hope_nicole ;P x goes up by 1....y goes down by -3 it's a fancy slope equation (y2-y1)/(x2-x1) 2 random points (-16 + 13)/(4-3) -3/1 so -3x when x = 0..y = -4 right? so y = -3x - 4

OpenStudy (anonymous):

lol, no i was to busy writing a book! besides i was never good at those, i could figure out the pattern(the up one down -3) but could never put it in an equation. @johnweldon1993

OpenStudy (johnweldon1993):

@Hope_nicole just messing with you anyways :) well now you know how :)

OpenStudy (anonymous):

yeah i know i know :) and yeah i wont remember that :P

OpenStudy (anonymous):

The equation a = 640s gives the relationship between s square miles and a acres. Pam owns 7.5 square miles of farmland. How many acres does she own

OpenStudy (johnweldon1993):

well if a...is 640 times how many square miles she owns...just plug that in a = 640(7.5) what do you get?

OpenStudy (anonymous):

A. 480 acres B. 11.72 acres C. 4,800 acres D. 1.17 acres

OpenStudy (johnweldon1993):

well heres a hint....640 is bigger than every number in that answer key except?

OpenStudy (anonymous):

4,800

OpenStudy (johnweldon1993):

exactly

OpenStudy (anonymous):

so it's c

OpenStudy (johnweldon1993):

correct

OpenStudy (anonymous):

The temperature on a particular day started at –13°F. It rose steadily by 3° each hour. The function y = –13 + 3x models the temperature, where x is the number of hours and y is the temperature. Graph the equation

OpenStudy (anonymous):

use slope intercept form, so basically flip flop the -13 and 3x to have y=3x-13 m is the slope, which in this case is 3 b is the y interept, which in this case is what?

OpenStudy (johnweldon1993):

well this is the slope intercept form y = mx + b m represents the slope of the line b represents the y-intercept the the line...that should be enough to graph your line

OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 3 2 2 3 1 4 0 A. f(x) = 4x B. f(x) = x – 4 C. f(x) = x + 4 D. f(x) = 4 – x

OpenStudy (anonymous):

again, try plugging in the data from the table into each equation till one satisfies all the points